w History What mass of NaNOs must be added to 450, g of waer to Mass Mole fraction propare a solution that is 0.0596 m in NaN07 What is the mole fractios of NaNO in the solation 3088
Answer #1

molar mass of NaNO2 = 69 g/mol
number of mole of NaNO2 = (given mass)/(molar mass)
= w/69

mass of water = 450 g
= 0.450 kg
molality = (number of mole of NaNO2)/(mass of water in kg)
0.0596 = (w/69)*(1/0.450)
w = 1.85 g

Answer : 1.85 g

molar mass of H2O = 18 g/mol
number of mole of NaNO2 = 1.85/69
= 0.027 mole
number of mole of H2O = 450/18
= 25 mole

mola fraction of NaNO2 = (number of mole of of NaNO2)/(total number of mole)
= 0.027/(25+0.027)
= (1.1*10^-3)

Answer : (1.1*10^-3)